3. Present Value. Compute the present value of $500 for the following combinations of annual percentage rates, years, and compounding periods: (a) APR-10%, n-5 years, m 365 (daily) (b) APR-10%, n 10 years, m-1 (annually) (c) APR-10%, n 20.25 years, m-4 (quarterly)
Answer #1

Present value is calculated using the excel function PV:

(a) =PV(10%/365,5*365,0,500,) = $303.2861

(b) =PV(10%,10,0,500,) = $192.7716

(c) =PV(10%/4,20.25*4,0,500,) = $67.6608